# Differential Gated Self-Attention

### v1

@@ Abstract
Transformers excel across a large variety of tasks but remain susceptible to corrupted inputs, since standard self-attention treats all query-key interactions uniformly. Inspired by lateral inhibition in biological neural circuits and building on the recent Differential Transformer’s use of two parallel softmax subtraction for noise cancellation, we propose Multihead Differential Gated Self-Attention (M-DGSA) that learns per-head input-dependent gating to dynamically suppress attention noise. Each head splits into excitatory and inhibitory branches whose dual softmax maps are fused by a sigmoid gate predicted from the token embedding, yielding a context-aware contrast enhancement. M-DGSA integrates seamlessly into existing Transformer stacks with minimal computational overhead. We evaluate on both vision and language benchmarks, demonstrating consistent robustness gains over vanilla Transformer, Vision Transformer, and Differential Transformer baselines. Our contributions are (i) a novel input-dependent gating mechanism for self-attention grounded in lateral inhibition, (ii) a principled synthesis of biological contrast-enhancement and self-attention theory, and (iii) comprehensive experiments demonstrating noise resilience and cross-domain applicability.
@@ 1 Introduction
Attention-based Transformers (Vaswani et al., 2017) have revolutionized machine learning, driving breakthroughs in language, vision, and beyond. However, their uniform weighting of all query-key scores makes them vulnerable to input corruption - sensor noise in images or spurious tokens in text - which can be propagated and even amplified degrading performance in real-world datasets.
Existing techniques for enhancing attention robustness or reducing its complexity, typically either apply global regularization (e.g. dropout, weight decay) or rework the attention computation itself via sparsity or low-rank factorization (Wang et al., 2020; Choromanski et al., 2020; Zaheer et al., 2020). More recently, the Differential Transformer (DIFF Transformer) (Ye et al., 2024) introduces a head-wise subtraction of two paired softmax maps to cancel common-mode noise, but its use of a single learned scalar limits its ability to adapt to low-level noise patterns that vary at the granularity of individual tokens.
In this work, we address the challenge of noise-cancellation in self-attention by introducing Multihead Differential Gated Self-Attention (M-DGSA), a lightweight module that learns an input-dependent inhibitory gating per-head to dynamically adapt to and suppress attention noise.
Inspired by lateral inhibition in biological neural circuits, where neurons suppress neighboring activity to sharpen responses (Hartline, 1940; Kuffler, 1953), M-DGSA splits each head into excitatory and inhibitory branches. Dual softmax maps are fused via a sigmoid gate predicted from the token embedding, enabling fine-grained contrast enhancement tailored to each query-key pair. M-DGSA integrates seamlessly into off-the-shelf Transformer-based stacks with negligible overhead.
We evaluate M-DGSA on both vision and language benchmarks - namely CIFAR-10, CIFAR-100, FashionMNIST, and SVHN for vision, Rotten Tomatoes, IMDB, AGNews, and 20Newsgroups for language - under both clean and noisy settings. Empirically, M-DGSA consistently improves accuracy and noise resilience across these domains, outperforming vanilla Transformer (Vaswani et al., 2017), Vision Transformer (ViT) (Dosovitskiy et al., 2020), and Differential Transformer (Ye et al., 2024) baselines. These results underscore the practical benefit of adaptive, input-dependent inhibition for attention-based architectures for robust performance across diverse downstream tasks.
@@
  • Adopting the lateral-inhibition principles of biological systems’ function into artificial attention mechanisms to provide an interpretable contrast-enhancement framework for self-attention,
@@ 2 Related Work
Building on the Transformer (Vaswani et al., 2017) architecture, many attention variants have been proposed to improve efficiency and robustness. Low-rank methods such as Linformer (Wang et al., 2020) and Performer (Choromanski et al., 2020) reduce complexity via kernel approximations, while sparsity-driven models like BigBird (Zaheer et al., 2020) leverage block-sparse patterns. In vision, the Vision Transformer (Dosovitskiy et al., 2020) demonstrates that pure self-attention over image patches, with an extra class token, can rival convolutional networks. ConViT (d’Ascoli et al., 2021) then introduces locality-sensitive gating to blend convolutional inductive biases with self-attention, enabling dynamic control over spatial focus. Recently, the Differential Transformer (Ye et al., 2024) introduces head-wise subtraction of parallel softmax maps to cancel common-mode noise, improve spectral balance and reduce attention collapse.
@@ 3.1 Scaled Dot-Product Attention
where $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d_{\mathrm{model}}\times d}$. The standard scaled dot-product attention then forms the weight matrix
@@
applying the softmax row-wise to normalize each query’s affinities, and the final output is
@@ 3.2 Differential Attention
$[Q_{1};Q_{2}]=X\,W_{Q},\quad[K_{1};K_{2}]=X\,W_{K},\quad V=X\,W_{V},$
with $W_{Q},W_{K}\in\mathbb{R}^{d_{\mathrm{model}}\times 2d^{\prime}}$ and $W_{V}\in\mathbb{R}^{d_{\mathrm{model}}\times 2d^{\prime}}$. Each stream then forms a standard scaled dot-product attention,
@@ 3.3 Lateral Inhibition
Lateral inhibition is a fundamental neural mechanism that enhances contrast by suppressing nearby activity. In rate-based models, the neuron’s response is characterized by
@@ 4 Method
We propose Differential Gated Self-Attention (M-DGSA), the first self-attention variant to embed an input-dependent lateral inhibition mechanism, combining the common-mode noise cancellation of differential attention with the adaptability of biological contrast-enhancement. Building on the Differential Transformer (Ye et al., 2024) implementation (dual-softmax streams, GroupNorm (Wu and He, 2018), and $\lambda_{init}$ scheme), we replace its subtraction scalar with a lightweight gating network that learns input-dependent and head-specific inhibitory weights.
4.1 Differential Gated Self-Attention
@@
Here we introduce a gating network, a per-input (token) $t$, per-head $h$ sigmoid gate
@@
1:$X\in\mathbb{R}^{N\times d_{model}}$, layer index $l$, weights $W_{Q},W_{K},W_{V},W_{g},W_{o}$, init $\lambda_{\mathrm{init}}$
2:$\lambda_{l}\leftarrow 0.8-0.6\exp(-0.3(l-1))$
3:$Q\leftarrow XW_{Q},\quad K\leftarrow XW_{K},\quad V\leftarrow XW_{V}$
4:Split $Q$ into $[Q^{+};Q^{-}]$, split $K$ into $[K^{+};K^{-}]$
5:$A^{+}\leftarrow\mathrm{softmax}(Q^{+}K^{+^{\top}}/{\sqrt{d}^{\prime}})$
6:$A^{-}\leftarrow\mathrm{softmax}(Q^{-}K^{-^{\top}}/\sqrt{d}^{\prime})$
7:$g\leftarrow\sigma(XW_{g}+b_{g})$
8:$A\leftarrow g\odot A^{+}-(1-g)\odot A^{-}$
9:$H\leftarrow(1-\lambda_{l})\mathrm{GroupNorm}(AV)$
10:$O\leftarrow$ reshape and concat heads of $H$
11:$Y\leftarrow Q+OW_{o}$
12:$Y\in\mathbb{R}^{N\times hd^{\prime}}$
@@
4.2 Multi-head Differential Gated Self-Attention
@@
$[\mathbf{Q^{+}};\mathbf{Q^{-}}],\,[\mathbf{K^{+}};\mathbf{K^{-}}]\;\in\;\mathbb{R}^{N\times 2h\times d^{\prime}},\quad\mathbf{V}\;\in\;\mathbb{R}^{N\times h\times 2d^{\prime}},$
and then reshape into $\{Q^{\pm}_{head_{i}},K^{\pm}_{head_{i}},V_{head_{i}}\}_{1}^{h}$, each of shape $N\times 2\times d^{\prime}$. Each head independently computes its fused map $A_{head_{i}}$ as in Eq. (1) and used to attend over $\mathbf{V}$, in the form of
@@
Concatenating $[head_{1},\dots,head_{i},\dots,head_{h}]$ along the embedding dimension and projecting with $W_{O}\in\mathbb{R}^{hd^{\prime}\times hd^{\prime}}$ and adding an optional residual connection from $[\mathbf{Q^{+}};\mathbf{Q^{-}}]$ to preserve information flow in our model. This completes the design of our multi-head attention block.
@@ 4.3 DGT and DGViT: Transformer & Vision Transformer Instantiations
Building on the ViT backbone, DGViT replaces each attention module with an M-DGSA layer, retaining the learnable class token in the attention computation and substitutes the GeLU activation (Hendrycks and Gimpel, 2016) in the feed-forward network with SwiGLU. It also integrates a residual connection inside the M-DGSA block to facilitate gradient flow, stabilize training, and boost model’s performance.
@@ 5 Experiments
In this section, we consider two important domains – vision and language classification - where we evaluate the performance of our differential gated self-attention mechanism. Detailed experimental setup and hyperparameter choices are provided in the supplementary materials.
All models are trained from scratch (no pre-trained weights), so our absolute accuracies may be lower than those reported in studies that fine-tune pre-trained networks. All experiments are implemented in PyTorch 2.7.0 and ran on a single machine equipped with an NVIDIA A10G GPU (24 GB VRAM).
@@ 5.1 Vision Classification
For vision classification, we leverage the built-in torchvision.datasets module, which offers a wide range of standardized benchmarks that integrate seamlessly with PyTorch for efficient data loading and preprocessing transformations. To evaluate robustness under challenging conditions, we inject noise into the training images across multiple distortion levels and noise types.
Datasets.
CIFAR-10 (Krizhevsky et al., 2009) contains 32×32 color images evenly distributed across 10 classes, and CIFAR-100 is its 100-class counterpart. Fashion-MNIST (Xiao et al., 2017) comprises 28×28 pixel grayscale Zalando article images spanning 10 fashion categories. The Street View House Numbers (SVHN) dataset (Netzer et al., 2011) consists of real-world house number images extracted from Google Street View.
@@
We compare our DGViT model against the original Vision Transformer (ViT) architecture, in which the feed-forward module expands the hidden dimension by a factor of two or four, applies a SwiGLU activation, and then projects back to the original size. To examine the impact of regularization, we insert one or two dropout layers interleaved between these projection and activation steps, thus assessing performance under different levels of stochastic masking throughout training.
Results.
Table 1 shows that the optimal configuration uses a 4×4$\times$4 × expansion followed by SwiGLU activation and a projection back to $d$, with two dropout layers. Relative to the baseline ViT, our DGViT model delivers competitive performance and marginal gains across all evaluated image-classification benchmarks. Notably, on CIFAR-10 we observe an absolute accuracy increase exceeding 2%. These findings indicate that expanding-and-shrinking the dimensionality around a SwiGLU activation, together with strategically placed dropout layers, promotes robust generalization across diverse visual domains and downstream tasks.
Table 1: Classification accuracy (mean $\pm$ std) on vision benchmarks. Averaged over 5 seeds.
@@ Results.
   ×$\times 2$$\times 2$ $77.51\pm 0.44$ $47.13\pm 0.26$ $93.42\pm 0.07$ $93.93\pm 0.34$
   ×$\times 4$$\times 1$ $77.77\pm 0.29$ $47.32\pm 0.38$ $93.32\pm 0.1$ $\mathbf{94.16\pm 0.49}$
   ×$\times 4$$\times 2$ $\mathbf{78.06\pm 0.41}$ $\mathbf{47.72\pm 0.35}$ $\mathbf{93.45\pm 0.13}$ $93.94\pm 0.18$

### v2

@@ Abstract
Transformers excel across a large variety of tasks but remain susceptible to corrupted inputs, since standard self‐attention treats all query-key interactions uniformly. Inspired by lateral inhibition in biological neural circuits and building on the recent Differential Transformer’s use of two parallel softmax subtraction for noise cancellation, we propose Multihead Differential Gated Self-Attention (M-DGSA) that learns per‐head input-dependent gating to dynamically suppress attention noise. Each head splits into excitatory and inhibitory branches whose dual softmax maps are fused by a sigmoid gate predicted from the token embedding, yielding a context-aware contrast enhancement. M-DGSA integrates seamlessly into existing Transformer stacks with minimal computational overhead. We evaluate on both vision and language benchmarks, demonstrating consistent robustness gains over vanilla Transformer, Vision Transformer, and Differential Transformer baselines. Our contributions are (i) a novel input-dependent gating mechanism for self‐attention grounded in lateral inhibition, (ii) a principled synthesis of biological contrast‐enhancement and self‐attention theory, and (iii) comprehensive experiments demonstrating noise resilience and cross-domain applicability.
@@ 1 Introduction
Attention‐based Transformers (Vaswani et al., 2017) have revolutionized machine learning, driving breakthroughs in language, vision, and beyond. However, their uniform weighting of all query-key scores makes them vulnerable to input corruption - sensor noise in images or spurious tokens in text - which can be propagated and even amplified degrading performance in real‐world datasets. Existing techniques for enhancing attention robustness or reducing its complexity, typically either apply global regularization (e.g. dropout, weight decay) or rework the attention computation itself via sparsity or low-rank factorization (Wang et al., 2020; Choromanski et al., 2020; Zaheer et al., 2020). More recently, the Differential Transformer (DIFF Transformer) (Ye et al., 2024) introduces a head-wise subtraction of two paired softmax maps to cancel common-mode noise, but its use of a single learned scalar limits its ability to adapt to low-level noise patterns that vary at the granularity of individual tokens.
In this work, we address the challenge of noise-cancellation in self‐attention by introducing Multihead Differential Gated Self‐Attention (M-DGSA), a lightweight module that learns an input-dependent inhibitory gating per‐head to dynamically adapt to and suppress attention noise.
Inspired by lateral inhibition in biological neural circuits, where neurons suppress neighboring activity to sharpen responses (Hartline, 1940; Kuffler, 1953), M-DGSA splits each head into excitatory and inhibitory branches. Dual softmax maps are fused via a sigmoid gate predicted from the token embedding, enabling fine‐grained contrast enhancement tailored to each query-key pair. M-DGSA integrates seamlessly into off‐the‐shelf Transformer-based stacks with negligible overhead.
We evaluate M‐DGSA on both vision and language benchmarks - namely CIFAR‐10, CIFAR-100, FashionMNIST, SVHN, and ImageNet-1k for vision, Rotten Tomatoes, IMDB, AGNews, 20Newsgroups, and MNLI for language - under both clean and noisy settings. Empirically, M-DGSA consistently improves accuracy and noise resilience across these domains, outperforming vanilla Transformer (Vaswani et al., 2017), Vision Transformer (ViT) (Dosovitskiy et al., 2020), and Differential Transformer (Ye et al., 2024) baselines. These results underscore the practical benefit of adaptive, input-dependent inhibition for attention‐based architectures for robust performance across diverse downstream tasks.
@@
  • Adopting the lateral‐inhibition principles of biological systems’ function into artificial attention mechanisms to provide an interpretable contrast-enhancement framework for self-attention,
@@ 2 Related Work
Building on the Transformer (Vaswani et al., 2017) architecture, many attention variants have been proposed to improve efficiency and robustness. Low-rank methods such as Linformer (Wang et al., 2020) and Performer (Choromanski et al., 2020) reduce complexity via kernel approximations, while sparsity-driven models like BigBird (Zaheer et al., 2020) leverage block-sparse patterns. In vision, the Vision Transformer (Dosovitskiy et al., 2020) demonstrates that pure self-attention over image patches, with an extra class token, can rival convolutional networks. ConViT (d’Ascoli et al., 2021) then introduces locality-sensitive gating to blend convolutional inductive biases with self-attention, enabling dynamic control over spatial focus. Recently, the Differential Transformer (Ye et al., 2024) introduces head-wise subtraction of parallel softmax maps to cancel common-mode noise, improve spectral balance and reduce attention collapse. Multi-Token Attention (Golovneva et al., 2025) performs a key–query convolution on the attention scores and applies a head-wise convolution across groups of attention heads to retrieve richer, reliable information from the input data.
@@ 3.1 Scaled Dot-Product Attention
where $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d_{\mathrm{model}}\times d}$. The standard scaled dot‐product attention then forms the weight matrix
@@
applying the softmax row‐wise to normalize each query’s affinities, and the final output is
@@ 3.2 Differential Attention
$Q_{1};Q_{2}]=X\,W_{Q},\quad[K_{1};K_{2}]=X\,W_{K},\quad V=X\,W_{V},$
with $W_{Q},W_{K}\in\mathbb{R}^{d_{\mathrm{model}}\times 2d^{\prime}}$ and $W_{V}\in\mathbb{R}^{d_{\mathrm{model}}\times 2d^{\prime}}$. Each stream then forms a standard scaled dot‐product attention,
@@ 3.3 Lateral Inhibition
Lateral inhibition is a fundamental neural mechanism that enhances contrast by suppressing nearby activity. In rate‐based models, the neuron’s response is characterized by
@@ 4 Method
We propose Differential Gated Self‐Attention (M‐DGSA), the first self‐attention variant to embed an input-dependent lateral inhibition mechanism, combining the common-mode noise cancellation of differential attention with the adaptability of biological contrast‐enhancement. Building on the Differential Transformer (Ye et al., 2024) implementation (dual‐softmax streams, GroupNorm (Wu and He, 2018), and $\lambda_{init}$ scheme), we replace its subtraction scalar with a lightweight gating network that learns input‐dependent and head‐specific inhibitory weights.
4.1 Differential Gated Self‐Attention
@@
Here we introduce a gating network, a per‐input (token) $t$, per‐head $h$ sigmoid gate
@@
[…]
4:Split $Q$ into $Q^{+};Q^{-}]$, split $K$ into $K^{+};K^{-}]$
[…]
@@
4.2 Multi-head Differential Gated Self‐Attention
@@
$\mathbf{Q^{+}};\mathbf{Q^{-}}],\,[\mathbf{K^{+}};\mathbf{K^{-}}]\;\in\;\mathbb{R}^{N\times 2h\times d^{\prime}},\quad\mathbf{V}\;\in\;\mathbb{R}^{N\times h\times 2d^{\prime}},$
and then reshape into $\{Q^{\pm}_{head_{i}},K^{\pm}_{head_{i}},V_{head_{i}}\}_{1}^{h}$, each of shape $N\times 2\times d^{\prime}$. Each head independently computes its fused map $A_{head_{i}}$ as in (1) and used to attend over $\mathbf{V}$, in the form of
@@
Concatenating $head_{1},\dots,head_{i},\dots,head_{h}]$ along the embedding dimension and projecting with $W_{O}\in\mathbb{R}^{hd^{\prime}\times hd^{\prime}}$ and adding an optional residual connection from $\mathbf{Q^{+}};\mathbf{Q^{-}}]$ to preserve information flow in our model. This completes the design of our multi‐head attention block.
@@ 4.3 DGT and DGViT: Transformer & Vision Transformer Instantiations
Building on the ViT backbone, DGViT replaces each attention module with an M-DGSA layer, retaining the learnable class token in the attention computation and substitutes the GeLU activation (Hendrycks and Gimpel, 2016) in the feed-forward network with SwiGLU. It also integrates an optional residual connection inside the M-DGSA block to facilitate gradient flow, stabilize training, and boost model’s performance.
@@ 5 Experiments
We evaluate the proposed Multihead Differential Gated Self-Attention (M-DGSA) across two representative domains: vision and natural language classification. The complete experimental setup, hyperparameter schedules, and dataset details are provided in Appendix A.
All models were trained from scratch (i.e., without pretrained weights) to isolate the contribution of the attention mechanism. As a result, absolute accuracies may be lower than prior work that fine-tunes pretrained backbones, however, this setting provides a controlled basis for cross-architecture comparison. Experiments were implemented in PyTorch 2.x and executed on a single machine equipped with an NVIDIA A10G (24 GB) for smaller datasets, and on servers with NVIDIA A100 (40 GB and 80 GB) GPUs for ImageNet-1k.
@@ 5.1 Vision Classification
For vision classification, we leverage the built-in torchvision.datasets module, which provides standardized benchmarks including, CIFAR-10 and CIFAR-100 (Krizhevsky et al., 2009), Fashion-MNIST (Xiao et al., 2017), and the Street View House Numbers (SVHN) dataset (Netzer et al., 2011). For large-scale evaluation, we additionally include ImageNet-1k (Deng et al., 2009). To assess robustness under challenging conditions, we further inject synthetic noise into the training images at multiple distortion levels.
@@
For the four smaller benchmarks, comparisons include the original Vision Transformer (ViT) and our Differential Gated ViT (DGViT). We also adapt the Differential Transformer (Ye et al., 2024) to the vision setting, denoted DViT. In DGViT and DViT, the feed-forward module expands the hidden dimension by a factor of two or four, applies a SwiGLU activation, and then projects back to the base size. To study the role of regularization, we interleave one or two dropout layers between these projection and activation steps, thereby assessing the effect of different levels of stochastic masking during training.
For ImageNet-1k, we compare the performance of ViT, DViT, and DGViT under two configurations: a $4\times$ expansion with two dropout layers, which serves as our default, and a $16/3\times$ expansion with two dropout layers, corresponding to the canonical setting of the Differential Transformer and providing a higher-capacity regime for comparison.
Results on small benchmarks.
Table 1 reports accuracies averaged over five seeds on CIFAR-10, CIFAR-100, FashionMNIST, and SVHN. DGViT consistently outperforms the baseline ViT, with the largest improvement on CIFAR-10 (+2.2%), highlighting the value of gated inhibition in low-resolution, noise-sensitive settings. On the other datasets, improvements are more modest (typically within 0.2-0.6%), yet they remain steady across configurations and never come at the cost of degraded accuracy. When compared directly to DViT, DGViT yields competitive or superior performance in all settings. On CIFAR-100, where DViT underperforms relative to ViT, DGViT closes this gap and establishes a clear advantage. On FashionMNIST, DGViT slightly surpasses DViT, demonstrating that input-dependent gating remains beneficial even in simpler domains. On SVHN, DGViT matches or marginally improves over DViT, confirming that the gating mechanism does not erode the gains provided by differential attention.
These narrow but steady margins across diverse datasets suggest that the contribution of the gating mechanism is not limited to specific domains but rather yields a systematic robustness benefit, complementing differential attention without introducing instability. In practice, this reliability is as important as large improvements on single datasets, as it demonstrates that the method scales consistently across different visual conditions.
Results on ImageNet-1k.
Table 2 summarizes performance at scale. Both DViT and DGViT substantially improve over the baseline ViT, with gains exceeding +6% absolute accuracy. Between the two differential variants, DGViT achieves consistent, though moderate, improvements over DViT for both expansion factors. These findings indicate that input-dependent gating provides an additional layer of robustness on top of differential attention, yielding the strongest overall performance while preserving computational efficiency.
Table 1: Classification accuracy (mean $\pm$ std) on small vision benchmarks. Averaged over 5 seeds.
@@
DViT
   proj. drop.
   $\times 2$ $\times 2$ $76.76\pm 0.28$ $45.49\pm 0.52$ $93.39\pm 0.10$ $93.79\pm 0.003$
   $\times 4$ $\times 1$ $76.40\pm 0.40$ $44.82\pm 0.41$ $93.32\pm 0.001$ $93.67\pm 0.001$
   $\times 4$ $\times 2$ $76.82\pm 0.42$ $45.55\pm 0.01$ $93.34\pm 0.001$ $93.80\pm 0.003$
   $\times 16/3$ $\times 2$ $77.16\pm 0.01$ $45.66\pm 0.26$ $93.34\pm 0.002$ $93.99\pm 0.001$
@@
   $\times 2$ $\times 2$ $77.51\pm 0.44$ $47.13\pm 0.26$ $42\pm 0.07 93.93$$93\pm 0.003$
   $\times 4$ $\times 1$ $77.77\pm 0.29$ $47.32\pm 0.38$ $93.32\pm 0.10$ $\mathbf{94.16\pm 0.004}$
   $\times 4$ $\times 2$ $78.06\pm 0.41$ $\mathbf{47.72\pm 0.35}$ $\mathbf{93.45\pm 0.13}$ $93.94\pm 0.002$
   $\times 16/3$ $\times 2$ $\mathbf{78.31\pm 0.01}$ $47.43\pm 0.03$ $\mathbf{93.45\pm 0.001}$ $93.89\pm 0.002$
Table 2: Classification accuracy (mean $\pm$ std) on ImageNet. Averaged over 5 seeds.
Model Accuracy (%)
ViT $59.94\pm 0.04$
DViT
   proj. drop.
   $\times 4$ $\times 2$ $65.75\pm 0.01$
   $\times 16/3$ $\times 2$ $66.66\pm 0.006$
DGViT (ours)
   proj. drop.
   $\times 4$ $\times 2$ $66.10\pm 0.004$
   $\times 16/3$ $\times 2$ $\mathbf{67.08\pm 0.003}$